Martha Manufacturing produces a single product that sells for $90per unit. Variable costs per unit equal $50. The company currently has total fixed costs of $72,000. Under the current cost and revenue structure Martha Manufacturing expects to sell 4,000 units. To improve short-term performance, management is considering a number of alternative actions. Evaluate each of the situations given below independently. Assume that, other than the information provided under the alternative actions, no other factors will change the revenue or cost structure.

1. Suppose management believes that a $20,000 increase in the monthly advertising expense will have a marked effect on sales. Sales must increase by how many units to justify this additional expenditure?

2.Suppose that management believes that a 10% reduction in the selling price will result in a 20% increase in number of units sold. Will this proposed reduction in selling price lead to an improvement in operating income? Show your calculations.

3.Suppose that management believes that a 10% increase in the selling price will result in a 15% reduction in number of units sold. Will this proposed reduction in selling price lead to an improvement in operating income? Show your calculations.
Answer #1

Contribution margin per unit = Sales price per unit – Variable cost per unit

= 90-50

= $40 per unit

Fixed costs and variable costs per unit remains constant within the relevant range

1.Increase in advertising expense = $20,000

Increase in units required = Increase in expense/CM per unit

= 20,000/40

= 500 units

2.Current operating Income = 40*4000-72000 = $88000

Operating income after change = (81 – 50)*4800 – 72000 = $76,800

No, will not lead to improvement

3.Operating income after change = (99-50)*3400 – 72000

= $94,600

Yes, will lead to improvement